OpenStudy (anonymous):

You have been asked to design a "ballistic spring system" to measure the speed of bullets. A spring whose spring constant k is is suspended from the ceiling. A block of mass M hangs from the spring. A bullet of mass m is fired vertically upward into the bottom of the block. The spring's maximum compression d is measured. Find an expression for the bullet's speed, Express your answer in terms of the variables m, M, k, d, and constant g.

OpenStudy (anonymous):

to begin with the spring was at rest even with the weight mg acting down so immediately there is a built-in restoring force of mg in the spring kinetic energy of the bullet is used to do work against gravity(-mgd) which is stored as +mgd and also to do work by spring force -kx =1/2kx^2 so kinetic====w.d by gravity+w.d by spring force

OpenStudy (mos1635):

bullet-block m*v=(M+m)*V =>V=..... spring-block(no bullet yet) M*g=k*d1 => d1=...... systemw energy after bullet Einitial=Efinal 1/2*(M+m)*V^2 + 1/2*k*d1^2= (M+m)*g*d + 1/2*k*(d1-d)^2